OpenStudy (anonymous):

what is the reciprocal of 6/5y?

OpenStudy (anonymous):

5y\6

OpenStudy (anonymous):

actually i think its supposed to be negative cause its the reciprocal

OpenStudy (anonymous):

yeah yeah right negative :)

OpenStudy (anonymous):

what about x/4 -x/6 = 0.5 ??

OpenStudy (anonymous):

give us a medal first ;)

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

make both dinominator the same \[6 * \frac{ x }{ 4} and 4* \frac{ x }{ 6 }\]

OpenStudy (anonymous):

next to my replies theres a tab called best response just click on it

OpenStudy (anonymous):